# A Multi-Fidelity Control Variate Approach for Policy Gradient Estimation

### v3

@@ 4 Multi-Fidelity Policy Gradients
Multi-fidelity policy gradient estimators via control variates. Because our objective is to optimize the policy performance in the high-fidelity environment $\mathcal{M}^{h}$, during each step of the policy gradient algorithm we must estimate $\nabla_{\theta}\mathbb{E}_{\tau^{h}\sim\mathcal{M}^{h}(\pi_{\theta})}[X_{\tau^{h}}^{\pi_{\theta}}]$ from a potentially limited number $N^{h}$ of sampled high-fidelity trajectories $\tau^{h}$. In data-scarce settings, existing policy gradient methods can face the challenge of high variance of the gradient estimates (Greensmith et al., 2004). We aim to reduce the estimation variance for the PGs. In this work, we assume that we may also sample a relatively large number $N^{l}\gg N^{h}$ of trajectories $\tau^{l}$ from the low-fidelity environment $\mathcal{M}^{l}$. We use these low-fidelity samples to construct a so-called control variate $X^{\pi_{\theta}}_{\tau^{l}}$—a correlated auxiliary random variable whose expected value ${\mu}^{l}\vcentcolon=\mathbb{E}_{\tau^{l}\sim\mathcal{M}^{l}(\pi_{\theta})}[X^{\pi_{\theta}}_{\tau^{l}}]$ is known. We then use the control variates technique (Nelson, 1987) to construct an unbiased, reduced-variance estimator for $\nabla_{\theta}\mathbb{E}_{\tau^{h}\sim\mathcal{M}^{h}(\pi_{\theta})}[X_{\tau^{h}}^{\pi_{\theta}}]$.
Specifically, we construct a new random variable $Z^{\pi_{\theta}}\vcentcolon=X_{\tau^{h}}^{\pi_{\theta}}+c(X_{\tau^{l}}^{\pi_{\theta}}-\mu^{l})$ with coefficient $c\in\mathbb{R}$. We estimate $\mu^{l}$ using the $N^{l}$ low-fidelity trajectory samples, and we provide a method to jointly sample correlated values $X_{\tau^{h}}^{\pi_{\theta}}$ and $X_{\tau^{l}}^{\pi_{\theta}}$ (described below). By construction, this new random variable has the property that $\mathbb{E}[Z^{\pi_{\theta}}]=\mathbb{E}[X_{\tau^{h}}^{\pi_{\theta}}]$. Furthermore, depending on the value of $c$, it may have a significantly reduced variance. In particular, by choosing
$c^{*}=-\frac{\textrm{Cov}(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})}{\textrm{Var}(X_{\tau^{l}}^{\pi_{\theta}})}=-\rho(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})\frac{\sqrt{\textrm{Var}(X_{\tau^{h}}^{\pi_{\theta}})}}{\sqrt{\textrm{Var}(X_{\tau^{l}}^{\pi_{\theta}})}}\implies\textrm{Var}(Z^{\pi_{\theta}})=(1-\rho^{2}(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}}))\textrm{Var}(X_{\tau^{h}}^{\pi_{\theta}}).$ (1)
Note that this choice of $c^{*}$ is a known closed-form solution minimizing the variance for $Z^{\pi_{\theta}}$ (Owen, 2013, Ch.8), and $\rho(\cdot,\cdot)$ is the Pearson correlation coefficient between the two random variables.
@@
In practice, we estimate $c^{*}$ using the sampled trajectories from $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$, reusing samples from previous gradient steps by maintaining moving-average statistics. Specifically, we maintain moving averages of the correlation $\hat{\rho}(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})$ and standard deviations $\sqrt{\widehat{\textrm{Var}}(X_{\tau^{h}}^{\pi_{\theta}})},\sqrt{\widehat{\textrm{Var}}(X_{\tau^{l}}^{\pi_{\theta}})}$. Upon each policy update, we compute
$\hat{c}^{*}=-\Bigl(\eta_{\text{ma}}$ $\hat{\rho}(X^{\pi_{\theta}}_{\tau^{h}},X^{\pi_{\theta}}_{\tau^{l}})_{\text{old}}+(1-\eta_{\text{ma}})\hat{\rho}(X^{\pi_{\theta}}_{\tau^{h}},X^{\pi_{\theta}}_{\tau^{l}})_{\text{new}}\Bigr)\frac{\eta_{\text{ma}}\sqrt{\widehat{\textrm{Var}}(X_{\tau^{h}}^{\pi_{\theta}})_{\text{old}}}+(1-\eta_{\text{ma}})\sqrt{\widehat{\textrm{Var}}(X_{\tau^{h}}^{\pi_{\theta}})_{\text{new}}}}{\eta_{\text{ma}}\sqrt{\widehat{\textrm{Var}}(X_{\tau^{l}}^{\pi_{\theta}})_{\text{old}}}+(1-\eta_{\text{ma}})\sqrt{\widehat{\textrm{Var}}(X_{\tau^{l}}^{\pi_{\theta}})_{\text{new}}}}.$ (2)
Here, the subscripts “old” and “new” denote the statistics carried over from the previous iteration and those estimated from the current batch, respectively. The hyperparameter $\eta_{\text{ma}}\in(0,1)$ controls the exponential moving-average weight.
@@
At every policy gradient step, the proposed MFPG framework thus proceeds as follows: 1) Use policy $\pi_{\theta}$ to sample $N^{h}$ correlated trajectories from $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$, as well as $N^{l}$ additional trajectories from $\mathcal{M}^{l}$. 2) Use the sampled trajectories to compute $\hat{\mu}^{l}$, $c$, and the correlated values of the random variables $X_{\tau^{h}}^{\pi_{\theta}}$ and $X_{\tau^{l}}^{\pi_{\theta}}$. 3) Compute the sampled values of $Z^{\pi_{\theta}}$. 4) Use the samples of $Z^{\pi_{\theta}}$ to compute an unbiased, reduced-variance estimate of the policy gradient $\nabla_{\theta}\mathbb{E}[Z^{\pi_{\theta}}]\approx\frac{1}{N^{h}}\sum_{\tau}\nabla_{\theta}\bigl[X^{\pi_{\theta}}_{\tau^{h}}+c(X^{\pi_{\theta}}_{\tau^{l}}-\hat{\mu}^{l})\bigr]$.
@@
Sampling correlated trajectories from the multi-fidelity environments. Note that for the random variables $X_{\tau^{h}}^{\pi_{\theta}}$ and $X_{\tau^{l}}^{\pi_{\theta}}$ to be correlated, they must share an underlying probability space $(\Omega,\mathcal{F},\text{I\kern-1.49994ptP})$. In other words, every outcome $\omega\in\Omega$ from this probability space should uniquely define a high-fidelity $\tau^{h}(\omega)$ and low-fidelity $\tau^{l}(\omega)$ trajectory, as well as the corresponding values of the random variables $X_{\tau^{h}}^{\pi_{\theta}}(\omega)$ and $X_{\tau^{l}}^{\pi_{\theta}}(\omega)$. Careful treatment of this joint probability space is necessary not only for conceptual completeness, but also to implement a mechanism that samples correlated trajectories from the distinct environments $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$ under the same policy $\pi_{\theta}$.
@@
$\textrm{TrajectoryList}\leftarrow\textrm{EmptyList}$
for $i\in\{1,2,\ldots,N^{h}\}$ do
 $\omega_{0}^{\pi}\ldots\omega_{T}^{\pi}\sim SamplePolicyOutcomes()$
 $s_{0}^{h}\sim\Delta_{s_{I}};s_{0}^{l}\leftarrow s_{0}^{h}$
  for $t\in\{0,\ldots,T-1\}$ do
     $a_{t}^{h}\leftarrow\pi_{\theta}(s_{t}^{h},\omega_{t}^{\pi_{\theta}})$
     $s_{t+1}^{h}\sim p^{h}(\cdot|s_{t}^{h},a_{t}^{h})$
     $r_{t}^{h}\sim R^{h}(s_{t}^{h},a_{t}^{h},s_{t+1}^{h})$
 for $t\in\{0,\ldots,T-1\}$ do
     $a_{t}^{l}\leftarrow\pi_{\theta}(s_{t}^{l},\omega_{t}^{\pi_{\theta}})$
     $s_{t+1}^{l}\sim p^{l}(\cdot|s_{t}^{l},a_{t}^{l})$
     $r_{t}^{l}\sim R^{l}(s_{t}^{l},a_{t}^{l},s_{t+1}^{l})$
 $\tau^{h}_{i}\leftarrow s_{0}^{h},a_{0}^{h},r_{0}^{h},\ldots,s_{T}^{h}$
 $\tau^{l}_{i}\leftarrow s_{0}^{l},a_{0}^{l},r_{0}^{l},\ldots,s_{T}^{l}$
 $\textrm{TrajectoryList.append}(\{\tau_{i}^{h},\tau_{i}^{l}\})$
return TrajectoryList

### v4

@@ 4 Multi-Fidelity Policy Gradients
Iteration notation.
We consider an iterative policy gradient procedure with parameters $\theta_{k}\in\mathbb{R}^{d}$ at iteration $k\in\mathbb{Z}^{+}$, and write $\pi_{\theta_{k}}(a\mid s)$ for the corresponding policy. When discussing a single iteration and a fixed parameter vector, all quantities are understood to be evaluated at that iteration, and we omit the subscript $k$ for readability, writing simply $\theta$ and $\pi_{\theta}$. All subsequent expressions that omit the iteration index should therefore be interpreted as taken at an arbitrary but fixed iteration $k$, and hence apply without loss of generality to every iteration of the algorithm.
Multi-fidelity policy gradient estimators via control variates.
@@
Because our objective is to optimize the policy performance in the high-fidelity environment $\mathcal{M}^{h}$, during each step of the policy gradient algorithm we must estimate $\nabla_{\theta}\mathbb{E}_{\tau^{h}\sim\mathcal{M}^{h}(\pi_{\theta})}[X_{\tau^{h}}^{\pi_{\theta}}]$ from a potentially limited number $N^{h}$ of sampled high-fidelity trajectories $\tau^{h}$. In data-scarce settings, existing policy gradient methods can face the challenge of high variance of the gradient estimates (Greensmith et al., 2004). We aim to reduce the estimation variance for the PGs. In this work, we assume that we may also sample a relatively large number $N^{l}\gg N^{h}$ of trajectories $\tau^{l}$ from the low-fidelity environment $\mathcal{M}^{l}$. We use these low-fidelity samples to construct a so-called control variate $X^{\pi_{\theta}}_{\tau^{l}}$—a correlated auxiliary random variable whose expected value ${\mu}^{l}\vcentcolon=\mathbb{E}_{\tau^{l}\sim\mathcal{M}^{l}(\pi_{\theta})}[X^{\pi_{\theta}}_{\tau^{l}}]$ is known. Note that $X^{\pi_{\theta}}_{\tau^{l}}$ is obtained by applying the same trajectory functional $X^{\pi_{\theta}}_{\tau}(\cdot)$ to trajectories $\tau^{l}$ sampled from the low-fidelity environment $\mathcal{M}^{l}(\pi_{\theta})$ under $\pi_{\theta}$. We then use the control variates technique (Nelson, 1987) to construct an unbiased, reduced-variance estimator for $\nabla_{\theta}\mathbb{E}_{\tau^{h}\sim\mathcal{M}^{h}(\pi_{\theta})}[X_{\tau^{h}}^{\pi_{\theta}}]$.
Specifically, we construct a new random variable $Z^{\pi_{\theta}}{\color[rgb]{0,0,0}(c)}\vcentcolon=X_{\tau^{h}}^{\pi_{\theta}}+c(X_{\tau^{l}}^{\pi_{\theta}}-\mu^{l})$ with coefficient $c\in\mathbb{R}$. The next lemma formalizes the unbiasedness and variance-reduction properties of this construction; its proof follows standard control variate arguments and is provided in Appendix˜B.
Lemma 1 (Unbiasedness and variance reduction of the multi-fidelity control variate estimator).
Let $X_{\tau^{h}}^{\pi_{\theta}}$ and $X_{\tau^{l}}^{\pi_{\theta}}$ be the high- and low-fidelity random variables defined above, and let ${\mu}^{l}\vcentcolon=\mathbb{E}_{\tau^{l}\sim\mathcal{M}^{l}(\pi_{\theta})}[X^{\pi_{\theta}}_{\tau^{l}}]$. For any coefficient $c\in\mathbb{R}$, define
$Z^{\pi_{\theta}}(c)\vcentcolon=X_{\tau^{h}}^{\pi_{\theta}}+c(X_{\tau^{l}}^{\pi_{\theta}}-\mu^{l}).$ (1)
Then, $Z^{\pi_{\theta}}(c)$ is unbiased with respect to the high-fidelity random variable: $\mathbb{E}[Z^{\pi_{\theta}}(c)]=\mathbb{E}[X_{\tau^{h}}^{\pi_{\theta}}]$.
Moreover, by choosing the coefficient
$c^{*}=-\frac{\textrm{Cov}(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})}{\textrm{Var}(X_{\tau^{l}}^{\pi_{\theta}})}=-\rho(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})\frac{\sqrt{\textrm{Var}(X_{\tau^{h}}^{\pi_{\theta}})}}{\sqrt{\textrm{Var}(X_{\tau^{l}}^{\pi_{\theta}})}}\implies\textrm{Var}\bigl(Z^{\pi_{\theta}}(c^{*})\bigr)=\bigl(1-\rho^{2}(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})\bigr)\textrm{Var}(X_{\tau^{h}}^{\pi_{\theta}}),$ (2)
where $\rho(\cdot,\cdot)$ is the Pearson correlation coefficient between the two random variables. In particular, whenever $\rho(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})\neq 0$, the multi-fidelity control variate estimator $Z^{\pi_{\theta}}(c^{*})$ has strictly smaller variance than $X_{\tau^{h}}^{\pi_{\theta}}$ alone.
@@
We estimate $\mu^{l}$ using the $N^{l}$ low-fidelity trajectory samples, and we provide a method to jointly sample correlated values $X_{\tau^{h}}^{\pi_{\theta}}$ and $X_{\tau^{l}}^{\pi_{\theta}}$ (described below). In the remainder of the paper, we take $Z^{\pi_{\theta}}\vcentcolon=Z^{\pi_{\theta}}(c^{*})$ as the ideal multi-fidelity random variable. In practice, $c^{*}$ is unknown and must be estimated from samples. We estimate $c^{*}$ using the sampled trajectories from $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$, reusing samples from previous gradient steps by maintaining moving-average statistics. Specifically, at iteration $k$, given the current policy $\pi_{\theta_{k}}$, we first compute batch estimates for correlation $\hat{\rho}^{\text{batch}}_{k}(X_{\tau^{h}}^{\pi_{\theta_{k}}},X_{\tau^{l}}^{\pi_{\theta_{k}}})$ and standard deviations $\sqrt{\widehat{\textrm{Var}}^{\text{batch}}_{k}(X_{\tau^{h}}^{\pi_{\theta_{k}}})},\sqrt{\widehat{\textrm{Var}}^{\text{batch}}_{k}(X_{\tau^{l}}^{\pi_{\theta_{k}}})}$ using newly sampled trajectories at iteration $k$. Let $\hat{\rho}_{k-1}(X_{\tau^{h}}^{\pi_{\theta_{k-1}}},X_{\tau^{l}}^{\pi_{\theta_{k-1}}})$, $\sqrt{\widehat{\textrm{Var}}_{k-1}(X_{\tau^{h}}^{\pi_{\theta_{k-1}}})},\text{and}\sqrt{\widehat{\textrm{Var}}_{k-1}(X_{\tau^{l}}^{\pi_{\theta_{k-1}}})}$ denote the running estimates from the previous iteration. With moving–average coefficient $\eta_{\text{ma}}\in(0,1)$, we update
$\hat{\rho}_{k}(X_{\tau^{h}}^{\pi_{\theta_{k}}},X_{\tau^{l}}^{\pi_{\theta_{k}}})$ $=\eta_{\text{ma}}\hat{\rho}_{k-1}(X_{\tau^{h}}^{\pi_{\theta_{k-1}}},X_{\tau^{l}}^{\pi_{\theta_{k-1}}})+(1-\eta_{\text{ma}})\,\hat{\rho}^{\text{batch}}_{k}(X_{\tau^{h}}^{\pi_{\theta_{k}}},X_{\tau^{l}}^{\pi_{\theta_{k}}}),$ (3)
$\sqrt{\widehat{\textrm{Var}}_{k}(X_{\tau^{h}}^{\pi_{\theta_{k}}})}$ $=\eta_{\text{ma}}\sqrt{\widehat{\textrm{Var}}_{k-1}(X_{\tau^{h}}^{\pi_{\theta_{k-1}}})}+(1-\eta_{\text{ma}})\,\sqrt{\widehat{\textrm{Var}}^{\text{batch}}_{k}(X_{\tau^{h}}^{\pi_{\theta_{k}}})},$
$\sqrt{\widehat{\textrm{Var}}_{k}(X_{\tau^{l}}^{\pi_{\theta_{k}}})}$ $=\eta_{\text{ma}}\sqrt{\widehat{\textrm{Var}}_{k-1}(X_{\tau^{l}}^{\pi_{\theta_{k-1}}})}+(1-\eta_{\text{ma}})\,\sqrt{\widehat{\textrm{Var}}^{\text{batch}}_{k}(X_{\tau^{l}}^{\pi_{\theta_{k}}})}.$
We then form an estimate of the optimal control–variate coefficient as
$\hat{c}_{k}^{*}\;=\;-\,\hat{\rho}_{k}(X_{\tau^{h}}^{\pi_{\theta_{k}}},X_{\tau^{l}}^{\pi_{\theta_{k}}})\,\frac{\sqrt{\widehat{\textrm{Var}}_{k}(X_{\tau^{h}}^{\pi_{\theta_{k}}})}}{\sqrt{\widehat{\textrm{Var}}_{k}(X_{\tau^{l}}^{\pi_{\theta_{k}}})}},$ (4)
and, when the iteration index is clear, we simply write $\hat{c}^{*}$.
@@
At every policy gradient step $k$, given the current policy $\pi_{\theta_{k}}$, the proposed MFPG framework thus proceeds as follows: 1) Use policy $\pi_{\theta_{k}}$ to sample $N^{h}$ correlated trajectories from $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$, as well as $N^{l}$ additional trajectories from $\mathcal{M}^{l}$. 2) Use the sampled trajectories to compute estimates $\hat{\mu}_{k}^{l}$, $\hat{c}^{*}_{k}$, and the correlated values of the random variables $X_{\tau^{h}}^{\pi_{\theta_{k}}}$ and $X_{\tau^{l}}^{\pi_{\theta_{k}}}$. 3) Compute the sampled values of $Z^{\pi_{\theta_{k}}}$. 4) Use the samples of $Z^{\pi_{\theta_{k}}}$ to compute an unbiased, reduced-variance estimate of the policy gradient ${\color[rgb]{0,0,0}\nabla_{\theta_{k}}\mathbb{E}[Z^{\pi_{\theta_{k}}}]}\approx\frac{1}{N^{h}}\sum_{\tau}{\color[rgb]{0,0,0}\nabla_{\theta_{k}}}\bigl[{\color[rgb]{0,0,0}X^{\pi_{\theta_{k}}}_{\tau^{h}}}+{\color[rgb]{0,0,0}\hat{c}^{*}_{k}}({\color[rgb]{0,0,0}X^{\pi_{\theta_{k}}}_{\tau^{l}}}-{\color[rgb]{0,0,0}\hat{\mu}_{k}^{l})}\bigr]$.
Sampling correlated trajectories from the multi-fidelity environments.
@@
Note that for the random variables $X_{\tau^{h}}^{\pi_{\theta}}$ and $X_{\tau^{l}}^{\pi_{\theta}}$ to be correlated, they must share an underlying probability space $(\Omega,\mathcal{F},\text{I\kern-1.49994ptP})$. In other words, every outcome $\omega\in\Omega$ from this probability space should uniquely define a high-fidelity $\tau^{h}(\omega)$ and low-fidelity $\tau^{l}(\omega)$ trajectory, as well as the corresponding values of the random variables $X_{\tau^{h}}^{\pi_{\theta}}(\omega)$ and $X_{\tau^{l}}^{\pi_{\theta}}(\omega)$. We emphasize that the global outcome $\omega$ is introduced purely as an analytical device. In practice, sampling and coupling the joint outcome $\omega$ across the high- and low-fidelity MDPs is infeasible, since the stochasticity in environment transitions and rewards is typically not directly controllable. Accordingly, our MFPG algorithm only couples the policy randomness and the initial-state randomness across fidelities, while treating transition and reward outcomes as independent, as we will describe below. A careful formulation of the joint probability space is nevertheless useful for conceptual clarity and for describing how to construct correlated trajectories in $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$ under the same policy $\pi_{\theta}$.
Informally, when sampling trajectories from $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$, there are six sources of stochasticity: the stochasticity introduced by the policy $\pi_{\theta}$, that introduced by the initial state distribution $\Delta_{s_{I}}$, the stochasticity introduced by the high-fidelity and low-fidelity transition dynamics $p^{h}$ and $p^{l}$, and finally the stochasticity introduced by the two reward functions $R^{h}$ and $R^{l}$. Note that the stochasticity introduced by $\pi_{\theta}$ is implemented by the agent, and is thus under the control of the algorithm in the sense that the same policy random outcome $\omega^{\pi}$ may be used to generate actions from $\pi_{\theta}$ in both $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$. We do so via a policy reparameterization trick (see below). Similarly, the algorithm may fix the initial states in the low-fidelity environment to match those observed from the high-fidelity samples. On the other hand, the transition dynamics and environment rewards are generated by independent sources of stochasticity in the different environments; we do not assume, nor require, any shared transition or reward outcome across fidelities.
We accordingly define the outcome set of the probability space as $\Omega=\Omega_{\Delta_{s_{I}}}\times\Omega_{\pi}\times\Omega_{p^{h}}\times\Omega_{p^{l}}\times\Omega_{R^{h}}\times\Omega_{R^{l}}$. Here, each outcome $\omega^{\Delta_{s_{I}}}\in\Omega_{\Delta_{s_{I}}}$ defines a particular shared initial state. Meanwhile, each policy outcome $\omega^{\pi}\in\Omega_{\pi}$ corresponds to a sequence $\omega^{\pi}=\omega^{\pi}_{1},\omega_{2}^{\pi},\ldots,\omega_{T}^{\pi}$ that dictates the random sequence of actions selected by the policy $\pi_{\theta}$ in both environments. In practice, this policy outcome sequence is realized as a sequence of per-timestep action-noise samples used to reparameterize the policy’s action sampling at each time step (see the discussion on policy reparameterization below). Conceptually, given the outcome $\omega_{t}^{\pi}$ at any timestep $t$ of a trajectory, the policy should deterministically output action $a^{h}_{t}=\pi_{\theta}(s_{t}^{h},\omega^{\pi})$ in the high-fidelity environment, and action $a^{l}_{t}=\pi_{\theta}(s_{t}^{l},\omega^{\pi})$ in the low-fidelity environment. Note that this does not necessarily imply that $a^{h}_{t}=a^{l}_{t}$, due to a potential difference in states $s^{h}_{t}$ and $s^{l}_{t}$ that stems from the different dynamics. Similarly, each outcome $\omega^{p_{h}}\in\Omega_{p^{h}}$ is a sequence $\omega^{p_{h}}=\omega_{0}^{p_{h}},\omega_{1}^{p_{h}},\ldots,\omega_{T}^{p_{h}}$ dictating the outcomes $s_{t+1}=p^{h}(s^{h}_{t},a^{h}_{t},\omega^{p_{h}})$ of the stochastic transitions in the high-fidelity environment. However, unlike the policy outcome sequence $\omega^{\pi}\in\Omega^{\pi}$, the transition outcome sequences $\omega^{p_{h}}\in\Omega^{p_{h}}$ and $\omega^{p_{l}}\in\Omega^{p_{l}}$, and the reward outcome sequences $\omega^{R_{h}}\in\Omega^{R_{h}}$ and $\omega^{R_{l}}\in\Omega^{R_{l}}$, are not shared between the high and low fidelity environments. Thus, only the initial-state outcomes $\omega^{\Delta_{s_{I}}}$ and the policy outcomes $\omega^{\pi}$ are coupled across fidelities in our algorithm, while the transition and reward outcomes $\omega^{p_{h}},\omega^{p_{l}},\omega^{R_{h}},\omega^{R_{l}}$ remain independent.
Formulating these separate outcome sets is conceptually helpful. However, practically, we only explicitly sample values for the policy outcomes $\omega_{0}^{\pi},\omega_{1}^{\pi},\ldots,\omega_{T}^{\pi}$ (action-noise samples) in our implementation, which we describe next.
@@
1
[…]
3
4for $i\in\{1,2,\ldots,N^{h}\}$ do
5
6 $\omega_{0}^{\pi}\ldots\omega_{T}^{\pi}\sim{\color[rgb]{0,0,0}SampleActionNoiseSequence()}$
7
8 $s_{0}^{h}\sim\Delta_{s_{I}};s_{0}^{l}\leftarrow s_{0}^{h}$
9
10 for $t\in\{0,\ldots,T-1\}$ do
    $a_{t}^{h}\leftarrow\pi_{\theta}(s_{t}^{h},\omega_{t}^{\pi_{\theta}})$ // reparameterization trick
11
12    $s_{t+1}^{h}\sim p^{h}(\cdot|s_{t}^{h},a_{t}^{h})$
13    $r_{t}^{h}\sim R^{h}(s_{t}^{h},a_{t}^{h},s_{t+1}^{h})$
14 for $t\in\{0,\ldots,T-1\}$ do
    $a_{t}^{l}\leftarrow\pi_{\theta}(s_{t}^{l},\omega_{t}^{\pi_{\theta}})$ // reparameterization trick
15
16    $s_{t+1}^{l}\sim p^{l}(\cdot|s_{t}^{l},a_{t}^{l})$
17    $r_{t}^{l}\sim R^{l}(s_{t}^{l},a_{t}^{l},s_{t+1}^{l})$
18
19 $\tau^{h}_{i}\leftarrow s_{0}^{h},a_{0}^{h},r_{0}^{h},\ldots,s_{T}^{h}$
20 $\tau^{l}_{i}\leftarrow s_{0}^{l},a_{0}^{l},r_{0}^{l},\ldots,s_{T}^{l}$
21 $\textrm{TrajectoryList.append}(\{\tau_{i}^{h},\tau_{i}^{l}\})$
22
23return TrajectoryList
24